OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form. ax^2 +bx +c=0 7 is the only solution. The equation is _=0

OpenStudy (anonymous):

If 7 is the only solution, and you want a quadratic function (a function that must have 2 solutions), 7 must be whats called a double root. There are a couple of ways to do this. One way (the most straightforward, not the easiest though) is to multiply this out: (x - 7)(x - 7) Another way (easier) is to know that if you have the roots of a polynomial, just add them, and multiply them, and that gives you the function. I want both roots to be 7. 7 + 7 = 14 7*7 = 49 so my answer is \[x^2-14x+49\]

OpenStudy (anonymous):

i guess i should be a little more specific, if you have the roots of a Quadratic polynomial, just add the root, and multiply them, and those numbers lead to your solution.

OpenStudy (anonymous):

More examples: I want a quadratic polynomial whose roots are 3 and -10 3+ (-10) = -7 3*(-10) = -30 answer is: \[x^2+7x-30\]